# KERMIT: Knowledge Graph Completion of Enhanced Relation Modeling with Inverse Transformation

### v2

@@
Abstract.
Knowledge graph completion (KGC) revolves around populating missing triples in a knowledge graph using available information. Text-based methods, which depend on textual descriptions of triples, often encounter difficulties when these descriptions lack sufficient information for accurate prediction—an issue inherent to the datasets and not easily resolved through modeling alone. To address this and ensure data consistency, we first use large language models (LLMs) to generate coherent descriptions, bridging the semantic gap between queries and answers. Secondly, we utilize inverse relations to create a symmetric graph, thereby providing augmented training samples for KGC. Additionally, we employ the label information inherent in knowledge graphs (KGs) to enhance the existing contrastive framework, making it fully supervised. These efforts have led to significant performance improvements on the WN18RR and FB15k-237 datasets. According to standard evaluation metrics, our approach achieves a 4.2% improvement in Hit@1 on WN18RR and a 3.4% improvement in Hit@3 on FB15k-237, demonstrating superior performance.
Knowledge graph completion (KGC), Large language models (LLMs), Supervised contrastive learning
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
Figure 1. An example of KGC from the FB15k-237 dataset in a text-based setting. Entities associated with their descriptions are connected by relational links.
1. Introduction
Knowledge graphs (KGs) have emerged as powerful tools for representing structured information, driving progress in areas like natural language processing, data mining, and artificial intelligence. KGs encode knowledge in triples $(h,r,t)$, where $h$ is the head entity, $t$ the tail entity, and $r$ the relationship, forming a rich semantic network. Their applications span recommendation systems (Guo et al., 2020), question answering (Fu et al., 2020), social network analysis (He et al., 2020b), and intelligent healthcare (MacLean, 2021; Gong et al., 2021). However, real-world KGs (e.g., Freebase (Bollacker et al., 2008), DBpedia (Auer et al., 2007), YAGO (Suchanek et al., 2007)) often suffer from incompleteness, missing vital relationships and entities (Dong et al., 2014). This issue drives the need for knowledge graph completion (KGC), which aims to fill these gaps by inferring missing information and enhancing the graph’s utility. Typically, a KGC model seeks to identify the most plausible tail entity $t$ for a given query $(h,r,?)$.
To address the KGC challenge, text-based methods (Yao et al., 2019; Wang et al., 2021b, 2022) have become increasingly sophisticated, with notable approaches representing triples through textual descriptions of the involved entities and relations. These methods leverage pre-trained language models (PLMs) for fine-tuning, encoding queries and answers as separate continuous vectors. A standout in this domain is SimKGC (Wang et al., 2022), which employs a bi-encoder contrastive learning framework to achieve state-of-the-art performance. Despite their advancements, most text-based KGC methods face two significant limitations due to data inconsistency: mismatched descriptions and pseudo-inverses.
@@ 1. Introduction
In addition, text-based methods often employ contrastive loss in a self-supervised manner, overlooking the label information inherent in a KG. In this paper, we introduce Knowledge graph completion of Enhanced Relation Modeling with Inverse Transformation (KERMIT) to address these challenges. To resolve the issue of ’mismatched description,’ we propose leveraging large language models (LLMs) as an external knowledge source to generate semantically coherent descriptions for KGC queries. Subsequently, we enhance the existing KG by meticulously crafting inverse relations, ensuring symmetry and coherence between queries and answers. It is important to emphasize that our method, as a fundamental data-level approach, can be applied to the majority of text-based methods, significantly improving their performance on KGC tasks. Lastly, we consider entities that satisfy the same query $(h,r,?)$ to belong to the same class, thereby enabling fully supervised contrastive learning (Khosla et al., 2020) to obtain more expressive representations. Empirical experiments validate the efficacy of our method, achieving state-of-the-art performance on the WN18RR (Toutanova et al., 2015) and FB15k-237 (Dettmers et al., 2018) datasets. In summary, our main contributions are as follows:
  • We propose to utilize LLMs to automatically address the ’mismatched descriptions’ and ’pseudo-inverse’ issues prevalent in text-based methods, with detailed discussions on prompting techniques and model selection.
@@
2. Related work
Text-based KGC methods. Recent advances in KGC have been primarily concentrated on scoring a triple $(h,r,t)$ based on the textual information associated with the entities and relation involved. Specifically, DKRL (Xie et al., 2016) encodes entity descriptions represented in bag-of-words format using a deep convolutional neural network (CNN). On the other hand, KG-BERT (Yao et al., 2019), StAR (Wang et al., 2021b) adopt two separate pre-trained language models, initialized with the same weights, to encode the queries and answers. SimKGC, as introduced in the work of (Wang et al., 2022), innovatively incorporates a contrastive learning framework (Chen et al., 2020) alongside several negative sampling methods. This integration results in a significant enhancement of text-based methods for KGC, highlighting the potential of contrastive learning in this context. Recently, CP-KGC (Zhang et al., 2023) utilizes LLMs to refine the existing textual descriptions of the entities, akin to our use of LLMs. However, CP-KGC focuses on improving description fluency, whereas our approach addresses the two specific issues we have identified.
Prompt engineering of LLMs. Prompt engineering has become essential for effective communication and interaction with LLMs (Marvin et al., 2023). Research (Kong et al., 2024; Shao et al., 2023a) suggests that assigning LLMs expert roles, a technique known as role-play prompting, can enhance their performance in specific domains. Large language models have shown a remarkable ability for in-context learning, meaning they can address various downstream tasks by conditioning on a prompt constructed from a few input-output examples (Ma et al., 2024). Additionally, instructing LLMs to provide extra explanations not only improves human interpretability but also increases the accuracy of their outputs (Jie et al., 2024). In our work, we combine these techniques to develop a well-crafted prompting template for guiding LLMs, addressing the challenges of ’mismatched descriptions’ and ’pseudo-inverse’.
@@ 2. Related work
Supervised contrastive learning. Contrastive learning has emerged as a powerful technique in unsupervised learning, significantly advancing representation learning by maximizing the similarity between augmented views of the same data point while minimizing it between different points (Jaiswal et al., 2021; Ben-Shaul et al., 2023). Seminal works like SimCLR (Chen et al., 2020) and MoCo (He et al., 2020a) have proven its effectiveness in learning visual representations without labels. The fundamental concept involves a contrastive loss function, such as InfoNCE (Chen et al., 2020; Yang et al., 2019), which promotes the learning of informative and discriminative features. To leverage label information effectively, (Khosla et al., 2020) extends the self-supervised batch contrastive approach to a fully-supervised setting, accommodating multiple positives per anchor. Recent research has further extended this method to various domains, including natural language processing (Zhao et al., 2023; Sidiropoulos and Kanoulas, 2024), showcasing its versatility and potential for improving downstream task performance. Nonetheless, this approach has not been widely adopted for graph data.
3. Preliminaries
@@
Text-based methods for KGC. To ensure this paper is self-contained, we will provide a general introduction to the principles of text-based methods. Typically, a text-based model is designed to learn a score function $f(\mathcal{D}(h),\mathcal{D}(r),\mathcal{D}(t))$ for a specific triple $(h,r,t)$ based on its textual description, where a higher score indicates greater plausibility. For simplicity, we denote the score function as $f(h,r,t)$. The majority of text-based approaches employ BERT (Devlin et al., 2019) to encode the textual information of a triple $(h,r,t)$ into semantic space. Two independently functioning BERT encoders with the same initial pre-trained weights are used to process the input $(h,r)$ and the output $t$. The embedding of the $(h,r)$ pair is noted as $e_{hr}$, while $e_{t}$ denotes the embedding of the tail entity. We will now examine the steps for generating $e_{hr}$ and $e_{t}$ embeddings:
@@
(1) $C_{hr}$ $=\text{BERT}_{\text{hr}}(X_{hr})$
(2) $C_{t}$ $=\text{BERT}_{\text{t}}(X_{t})$
@@
(3) $X_{hr}$ $=\left[x_{\text{CLS}},H_{\text{desc}},x_{\text{SEP}},R_{\text{desc}},x_{\text{SEP}}\right]$
(4) $X_{t}$ $=\left[x_{\text{CLS}},T_{\text{desc}},x_{\text{SEP}}\right]$
where $x_{\text{CLS}}$ and $x_{\text{SEP}}$ are special tokens introduced by (Devlin et al., 2019). $C_{hr}$ and $C_{t}$ are the corresponding representation vectors.
@@
(5) $e_{hr}$ $=\text{Pool}(C_{hr})$
(6) $e_{t}$ $=\text{Pool}(C_{t})$
@@
(7) $\cos(e_{hr},e_{t})=\frac{e_{hr}\cdot e_{t}}{||e_{hr}||\cdot||e_{t}||}$
Finally, following the idea of contrastive learning (Jaiswal et al., 2021; Ben-Shaul et al., 2023) for acquiring expressive representations, InfoNCE loss, augmented with an additive margin (Chen et al., 2020; Yang et al., 2019), is employed to construct the training objective:
(8) $\mathcal{L}=-\log\frac{e^{(f(h,r,t)-\gamma)/\tau}}{e^{(f(h,r,t)-\gamma)/\tau}+\sum_{t^{\prime}\in\mathcal{N}^{-}}{e^{(f(h,r,t^{\prime})-\gamma)/\tau}}}$
@@
(9) $\hat{t}=\mathop{\arg\max}\limits_{t}\cos(e_{hr},e_{t}),t\in\mathcal{E}$
Prompting LLMs. Basically, we utilize LLMs to retrieve responses with the prompting technique (Liu et al., 2023) in a question-and-answer process. To do this, the initial step involves the design of a prompting template denoted as P, which incorporates a placeholder {X} intended for accommodating the input $\mathbf{x}$. Subsequently, a specific input $\mathbf{x}$ is inserted into this placeholder, thereby creating the prompt that serves as the direct input to the LLMs. Finally, we collect the generated output $\mathbf{y}$ produced by the LLMs. This procedural sequence can be formally articulated as follows:
(10) $\mathbf{y}=\text{LLM}(\mathbf{x}|\textbf{P})$
Table 1. The prompting template used to generate the predictive description of the tail entity includes placeholders for the head entity’s name (h_name) and its textual description (h_desc), the relation name (r_name), and a few-shot example of triples (triple_examples). These placeholders will be replaced when dealing with an incomplete triple $(h,r,?)$.
You are a learned knowledge creator. I have an entity ’{h_name}’ and a relation ’{r_name}’. Here is the description of the entity:
@@
There are several examples for you in which two entities satisfy the relation ’{r_name}’. Each example is in the form of a triple (h, r, t), meaning that the relationshiop between the entities ’h’ and ’t’ is ’r’. For instance, the fact that ’Bob is the father of Tom’ can be expressed by the triple (Bob, father of, Tom). Each example takes one single line:
@@
Your task is to find an entity such that the relationship between ’{h_name}’ and the entity is ’{r_name}’. First, list at least five potential entities that might satisfy the requirement. Then, generate a broad description of the common properties shared by the entities. Your response must follow the JSON format below:
@@
The description is limited to 30 words and no less than 10 words. The explanation should detail why you choose the entities based on the given entity ’{h_name}’, relation ’{r_name}’ and triple examples.
Table 2. The prompting template used to generate inverse relations includes placeholders the relation name (r_name), and a few-shot example of triples (triple_examples). These placeholders will be replaced when dealing with a specific relation $r$.
You are a learned knowledge creator. I have a relation ’{r_name}’. There are several examples for you in which two entities satisfy the relation ’{r_name}’. Each example is in the form of a triple (h, r, t), meaning that the relationship between the entities ’h’ and ’t’ is ’r’. For instance, the fact that ’Bob is the father of Tom’ can be expressed by the triple (Bob, father of, Tom). Each example takes one single line:
@@
Your task is to find an inverse (r’) of the relation ’{r_name}’ such that the form (t, r’, h) holds true for all the given triples. Generate a potential relation that might satisfy the requirement. Your response must follow the JSON format below:
@@
The explanation should detail why you choose the inverse based on the given relation ’{r_name}’ and triple examples.
Figure 2. An overview of our approach, KERMIT. For a given query $(h,r,?)$, KERMIT begins by sampling a subset of positive entities $\{t|t\in\mathcal{E},(h,r,t)\in\mathcal{T}\}$ and a subset of negative entities $\{t^{\prime}|t^{\prime}\in\mathcal{E},(h,r,t^{\prime})\notin\mathcal{T}\}$. The query and potential answer entities are then converted into token sequences, followed by two separate BERT encoders to generate embeddings. Finally, a supervised contrastive loss is applied to optimize the encoders.
4. Methodology
This section first proposes our method, namely ’predictive description’, designed to tackle the challenges of ’mismatched description’ and ’pseudo-inverse’ with the aid of LLMs. Following this, we will provide an detailed explanation of our model’s architecture and training strategy, incorporating a supervised contrastive loss built upon the inherent label information of a KG.
4.1. Predictive description via LLMs
Under the formulation in Section 3, the KGC task here is considered to be an entity prediction problem (Malaviya et al., 2020). From this point of view, the textual description plays an important role in featuring an entity.
Although encoding a triple through its textual information yields good performance for KGC tasks, relying solely on descriptions of the head entity and relation usually fails to accurately predict the tail. The issue arises when the description of a query $(h,r)$ lacks coherence with the tail $t$, resulting in suboptimal predictions. Thereby, we propose to use LLMs, as they can serve as an implicit database for vast open-domain knowledge (Shao et al., 2023b; Li et al., 2022), to generate predictive information aimed at the tail entity with a well-designed prompting template $\textbf{P}_{\text{desc}}$:
(11) $t_{\text{pred}}=\text{LLM}(g_{k}(h,r)|\textbf{P}_{\text{desc}})$
The prompting template we use, as part of the input to LLMs, is shown in Table 1. The function $g_{k}(\cdot)$ is a pre-processing step that fetches textual descriptions and retrieves $k$ few-shot triple examples sharing the same relation $r$. We fix the number of few-shot examples $k$ to $5$ during all experiments. In this prompt, we utilize role-playing (Kong et al., 2024; Shao et al., 2023a) and few-shot engineering techniques (Ma et al., 2024) to instruct the LLM to generate a list of potential tail entities that could complete the query $(h,r,?)$, along with general descriptions of these entities. Additionally, an explanation from the LLM is requested to ensure accuracy (Jie et al., 2024). Finally, we employ simple concatenation to obtain the predictive description of the query: ’{entities}: {description}’.
@@ 4.1. Predictive description via LLMs
(12) $h_{\text{pred}}=\text{LLM}(g_{k}(t,r^{\prime})|\textbf{P}_{\text{desc}})$
LLMs offer a cost-effective and efficient mechanism for obtaining textual information, obviating the necessity for human intervention. We also notice that some recent methods solve entity prediction problems by involving LLMs as external knowledge engines, e.g. Prophet (Shao et al., 2023b), a knowledge-based visual question answering (VQA) model, teaches GPT-3 to answer visual questions from given candidates by showing examples selected by a vanilla VQA model to it, which is a certain case of in-context learning (Dong et al., 2022). The task of KGC might also be handled through the combination of a vanilla KGC model and an LLM. We leave the investigation of this idea for future work.
Addressing the challenge of ’pseudo-inverse’ in KGC, it is important to note that each node can act both as a head and a tail. Considering this duality, we propose converting the relations to their inverse forms to maintain semantic coherence. Similarly, we utilize LLMs to identify these inverse relations, employing the prompting template displayed in Table 2:
(13) $r^{\prime}=\text{LLM}(g_{k}(r)|\textbf{P}_{\text{rel}})$
@@
4.2. Model architecture

### v3

# mode = titleKERMIT: Knowledge Graph Completion of Enhanced Relation Modeling with Inverse Transformation

@@
Abstract
Knowledge graph completion (KGC) revolves around populating missing triples in a knowledge graph using available information. Text-based methods, which depend on textual descriptions of triples, often encounter difficulties when these descriptions lack sufficient information for accurate prediction, an issue inherent to the datasets and not easily resolved through modeling alone. To address this and ensure data consistency, we first use large language models (LLMs) to generate coherent descriptions, bridging the semantic gap between queries and answers. Secondly, we utilize inverse relations to create a symmetric graph, thereby providing augmented training samples for KGC. Additionally, we employ the label information inherent in knowledge graphs (KGs) to enhance the existing contrastive framework, making it fully supervised. These efforts have led to significant performance improvements on the WN18RR, FB15k-237 and UMLS datasets. According to standard evaluation metrics, our approach achieves a 3.0% improvement in Hit@1 on WN18RR and a 12.1% improvement in Hit@3 on UMLS, demonstrating superior performance.
keywords:
@@ Abstract
1 Introduction
Knowledge graphs (KGs) have emerged as powerful tools for representing structured information, driving progress in areas like natural language processing, data mining, and artificial intelligence. KGs encode knowledge in triples $(h,r,t)$, where $h$ is the head entity, $t$ the tail entity, and $r$ the relationship, forming a rich semantic network. Their applications span recommendation systems [1], question answering [2], social network analysis [3], and cyber security [4]. However, real-world KGs (e.g., Freebase [5], DBpedia [6], YAGO [7]) often suffer from incompleteness, missing vital relationships and entities [8]. This issue drives the need for knowledge graph completion (KGC), which aims to fill these gaps by inferring missing information and enhancing the graph’s utility. Typically, a KGC model seeks to identify the most plausible tail entity $t$ for a given query $(h,r,?)$.
Figure 1: An example of KGC from the FB15k-237 dataset in a text-based setting. Entities associated with their descriptions are connected by relational links.
To address the KGC challenge, text-based methods [9, 10, 11] have become increasingly sophisticated, with notable approaches representing triples through textual descriptions of the involved entities and relations. These methods leverage pre-trained language models (PLMs) for fine-tuning, encoding queries and answers as separate continuous vectors. A standout in this domain is SimKGC [11], which employs a bi-encoder contrastive learning framework to achieve state-of-the-art performance. Despite their advancements, most text-based KGC methods face two significant limitations due to data inconsistency: mismatched descriptions and pseudo-inverses.
[…]
In addition, text-based methods often employ contrastive loss in a self-supervised manner, overlooking the label information inherent in a KG. In this paper, we introduce Knowledge graph completion of Enhanced Relation Modeling with Inverse Transformation (KERMIT) to address these challenges. To resolve the issue of "mismatched description", we propose leveraging large language models (LLMs) as an external knowledge source to generate semantically coherent descriptions for KGC queries. Subsequently, we enhance the existing KG by meticulously crafting inverse relations, ensuring symmetry and coherence between queries and answers. It is important to emphasize that our method, as a fundamental data-level approach, can be applied to the majority of text-based methods, significantly improving their performance on KGC tasks. Lastly, we consider entities that satisfy the same query $(h,r,?)$ to belong to the same class, thereby enabling fully supervised contrastive learning [13] to obtain more expressive representations. Empirical experiments validate the efficacy of our method, achieving state-of-the-art performance on the WN18RR [14] and FB15k-237 [12] datasets. In summary, our main contributions are as follows:
  • We propose to utilize LLMs to automatically address the "mismatched descriptions" and "pseudo-inverse" issues prevalent in text-based methods, with detailed discussions on prompting techniques and model selection.
@@ 1 Introduction
2 Related work
Knowledge graph completion. The task of KGC has traditionally centered on designing effective scoring mechanisms to assess the plausibility of triples in low-dimensional embedding spaces. A foundational approach in this area is knowledge graph embedding (KGE), commonly known as embedding-based methods. Representative techniques include TransE [15], DistMult [16], RotatE [17], ConvE [12], and TuckER [18]. Despite their success in KGC performance, these approaches are often opaque, functioning as black boxes and making them difficult for humans to understand. Alternatively, rule-based methods such as NeuralLP [19], DRUM [20], RNNLogic [21], and LERP [22] focus on learning first-order logical Horn clauses for KGC. Despite their clear interpretability, rule-based methods often struggle with scalability to large graphs and suffer from inefficiencies in training.
Text-based KGC methods. Recent developments in KGC have emphasized scoring a triple $(h,r,t)$ using textual information from the entities and relations involved. DKRL [23] leverages a deep convolutional neural network (CNN) to encode entity descriptions in a bag-of-words format. MTL-KGC [24] adopts a multi-task learning framework for text-based KGC tasks. Meanwhile, KG-BERT [9] and StAR [10] utilize two separate pre-trained language models with identical initial weights to encode queries and answers. SimKGC [11] introduces a contrastive learning framework [25] combined with advanced negative sampling methods, which is further improved by GHN [26] through the generation of semantic-enhanced negatives, resulting in notable advancements for text-based KGC. Recently, large language models (LLMs) have emerged as powerful tools, leading to methods like CP-KGC [27] and KICGPT [28] that leverage LLMs to advance KGC.
Prompt engineering of LLMs. Effective interaction with LLMs increasingly relies on prompt engineering [29]. Role-play prompting, a strategy where LLMs are assigned specific expert roles, has been shown to enhance domain-specific performance [30, 31]. Furthermore, LLMs exhibit exceptional in-context learning capabilities, handling various downstream tasks by conditioning on prompts consisting of a few illustrative input-output examples [32]. Encouraging LLMs to generate additional explanations improves both the interpretability of their outputs and their overall accuracy [33]. In this work, we employ these techniques to develop a tailored prompting template that effectively addresses challenges such as "mismatched descriptions" and "pseudo-inverse".
@@ 2 Related work
Supervised contrastive learning. Contrastive learning has become a cornerstone of unsupervised learning, significantly advancing representation learning by encouraging augmented views of the same data point to be similar, while pushing apart representations of different data points [34, 35]. Foundational works like SimCLR [25] and MoCo [36] have validated its potential for learning effective visual representations without requiring labels. At its heart lies the contrastive loss, such as InfoNCE [25, 37], which promotes learning features that are both informative and discriminative. The work of [13] extends this concept to a fully supervised batch contrastive setting, effectively leveraging label information by accommodating multiple positives per anchor. More recently, contrastive learning has been successfully extended to other domains, such as natural language processing [38, 39], illustrating its adaptability and potential. Despite these advances, its application to graph data remains underexplored.
3 Preliminaries
@@
Text-based methods for KGC. To ensure this paper is self-contained, we will provide a general introduction to the principles of text-based methods. Typically, a text-based model is designed to learn a score function $f(\mathcal{D}(h),\mathcal{D}(r),\mathcal{D}(t))$ for a specific triple $(h,r,t)$ based on its textual description, where a higher score indicates greater plausibility. For simplicity, we denote the score function as $f(h,r,t)$. The majority of text-based approaches employ BERT [40] to encode the textual information of a triple $(h,r,t)$ into semantic space. Two independently functioning BERT encoders with the same initial pre-trained weights are used to process the input $(h,r)$ and the output $t$. The embedding of the $(h,r)$ pair is noted as $e_{hr}$, while $e_{t}$ denotes the embedding of the tail entity. We will now examine the steps for generating $e_{hr}$ and $e_{t}$ embeddings:
@@
$C_{hr}$ $=\text{BERT}_{\text{hr}}(X_{hr}),~{}~{}C_{t}=\text{BERT}_{\text{t}}(X_{t})$ (1)
@@
$X_{hr}$ $=\left[x_{\text{CLS}},H_{\text{desc}},x_{\text{SEP}},R_{\text{desc}},x_{\text{SEP}}\right]$ (2)
$X_{t}$ $=\left[x_{\text{CLS}},T_{\text{desc}},x_{\text{SEP}}\right]$ (3)
where $x_{\text{CLS}}$ and $x_{\text{SEP}}$ are special tokens introduced by [40]. $C_{hr}$ and $C_{t}$ are the corresponding representation vectors.
@@
$e_{hr}$ $=\text{Pool}(C_{hr}),~{}~{}e_{t}=\text{Pool}(C_{t})$ (4)
@@
$\cos(e_{hr},e_{t})=\frac{e_{hr}\cdot e_{t}}{||e_{hr}||\cdot||e_{t}||}$ (5)
Finally, following the idea of contrastive learning [34, 35] for acquiring expressive representations, InfoNCE loss, augmented with an additive margin [25, 37], is employed to construct the training objective:
$\mathcal{L}=-\log\frac{e^{(f(h,r,t)-\gamma)/\tau}}{e^{(f(h,r,t)-\gamma)/\tau}+\sum_{t^{\prime}\in\mathcal{N}^{-}}{e^{(f(h,r,t^{\prime})-\gamma)/\tau}}}$ (6)
@@
$\hat{t}=\mathop{\arg\max}\limits_{t}\cos(e_{hr},e_{t}),t\in\mathcal{E}$ (7)
Prompting LLMs. Basically, we utilize LLMs to retrieve responses with the prompting technique [41] in a question-and-answer process. To do this, the initial step involves the design of a prompting template denoted as P, which incorporates a placeholder {X} intended for accommodating the input $\mathbf{x}$. Subsequently, a specific input $\mathbf{x}$ is inserted into this placeholder, thereby creating the prompt that serves as the direct input to the LLMs. Finally, we collect the generated output $\mathbf{y}$ produced by the LLMs. This procedural sequence can be formally articulated as follows:
$\mathbf{y}=\text{LLM}(\mathbf{x}|\textbf{P})$ (8)
Table 1: The prompting template used to generate the predictive description of the tail entity includes placeholders for the head entity’s name (h_name) and its textual description (h_desc), the relation name (r_name), and a few-shot example of triples (triple_examples). These placeholders will be replaced when dealing with an incomplete triple $(h,r,?)$.
You are a learned knowledge creator. I have an entity "{h_name}" and a relation "{r_name}". Here is the description of the entity:
@@
There are several examples for you in which two entities satisfy the relation "{r_name}". Each example is in the form of a triple (h, r, t), meaning that the relationship between the entities "h" and "t" is "r". For instance, the fact that "Bob is the father of Tom" can be expressed by the triple (Bob, father of, Tom). Each example takes one single line:
@@
Your task is to find an entity such that the relationship between "{h_name}" and the entity is "{r_name}". First, list at least five potential entities that might satisfy the requirement. Then, generate a broad description of the common properties shared by the entities. Your response must follow the JSON format below:
@@
The description is limited to 30 words and no less than 10 words. The explanation should detail why you choose the entities based on the given entity "{h_name}", relation "{r_name}" and triple examples.
4 Methodology
This section first proposes our method, namely "predictive description", designed to tackle the challenges of "mismatched description" and "pseudo-inverse" with the aid of LLMs. Following this, we will provide an detailed explanation of our model’s architecture and training strategy, incorporating a supervised contrastive loss built upon the inherent label information of a KG.
4.1 Predictive description via LLMs
Under the formulation in Section 3, the KGC task here is considered to be an entity prediction problem [42]. From this point of view, the textual description plays an important role in featuring an entity.
Although encoding a triple through its textual information yields good performance for KGC tasks, relying solely on descriptions of the head entity and relation usually fails to accurately predict the tail. The issue arises when the description of a query $(h,r)$ lacks coherence with the tail $t$, resulting in suboptimal predictions. Thereby, we propose to use LLMs, as they can serve as an implicit database for vast open-domain knowledge [43, 44], to generate predictive information aimed at the tail entity with a well-designed prompting template $\textbf{P}_{\text{desc}}$:
$t_{\text{pred}}=\text{LLM}(g_{k}(h,r)|\textbf{P}_{\text{desc}})$ (9)
Table 2: The prompting template used to generate inverse relations includes placeholders the relation name (r_name), and a few-shot example of triples (triple_examples). These placeholders will be replaced when dealing with a specific relation $r$.
You are a learned knowledge creator. I have a relation "{r_name}". There are several examples for you in which two entities satisfy the relation "{r_name}". Each example is in the form of a triple (h, r, t), meaning that the relationship between the entities "h" and "t" is "r". For instance, the fact that "Bob is the father of Tom" can be expressed by the triple (Bob, father of, Tom). Each example takes one single line:
@@ 4.1 Predictive description via LLMs
Your task is to find an inverse (r’) of the relation "{r_name}" such that the form (t, r’, h) holds true for all the given triples. Generate a potential relation that might satisfy the requirement. Your response must follow the JSON format below:
@@
The explanation should detail why you choose the inverse based on the given relation "{r_name}" and triple examples.
The prompting template we use, as part of the input to LLMs, is shown in Table 1. The function $g_{k}(\cdot)$ is a pre-processing step that fetches textual descriptions and retrieves $k$ few-shot triple examples sharing the same relation $r$. We fix the number of few-shot examples $k$ to $5$ during all experiments. In this prompt, we utilize role-playing [30, 31] and few-shot engineering techniques [32] to instruct the LLM to generate a list of potential tail entities that could complete the query $(h,r,?)$, along with general descriptions of these entities. Additionally, an explanation from the LLM is requested to ensure accuracy [33]. Finally, we employ simple concatenation to obtain the predictive description of the query: "{entities}: {description}".
@@
$h_{\text{pred}}=\text{LLM}(g_{k}(t,r^{\prime})|\textbf{P}_{\text{desc}})$ (10)
LLMs offer a cost-effective and efficient mechanism for obtaining textual information, obviating the necessity for human intervention. We also notice that some recent methods solve entity prediction problems by involving LLMs as external knowledge engines, e.g. Prophet [43], a knowledge-based visual question answering (VQA) model, teaches GPT-3 to answer visual questions from given candidates by showing examples selected by a vanilla VQA model to it, which is a certain case of in-context learning [45]. The task of KGC might also be handled through the combination of a vanilla KGC model and an LLM. We leave the investigation of this idea for future work.
Addressing the challenge of "pseudo-inverse" in KGC, it is important to note that each node can act both as a head and a tail. Considering this duality, we propose converting the relations to their inverse forms to maintain semantic coherence. Similarly, we utilize LLMs to identify these inverse relations, employing the prompting template displayed in Table 2:
$r^{\prime}=\text{LLM}(g_{k}(r)|\textbf{P}_{\text{rel}})$ (11)
@@
Figure 2: An overview of KERMIT during training. For a given query $(h,r,?)$, KERMIT begins by sampling a subset of positive entities $\{t|t\in\mathcal{E},(h,r,t)\in\mathcal{T}\}$ and a subset of negative entities $\{t^{\prime}|t^{\prime}\in\mathcal{E},(h,r,t^{\prime})\notin\mathcal{T}\}$. The query and potential answer entities are then converted into token sequences, followed by two separate BERT encoders to generate embeddings. Finally, a supervised contrastive loss is applied to optimize the encoders.
4.2 Model architecture